# Ensuring Force Safety in Vision-Guided Robotic Manipulation via Implicit Tactile Calibration

### v1

@@ Abstract
In dynamic environments, robots often encounter constrained movement trajectories when manipulating objects with specific properties, such as doors. Therefore, applying the appropriate force is crucial to prevent damage to both the robots and the objects. However, current vision-guided robot state generation methods often falter in this regard, as they lack the integration of tactile perception. To tackle this issue, this paper introduces a novel state diffusion framework termed SafeDiff. It generates a prospective state sequence from the current robot state and visual context observation while incorporating real-time tactile feedback to refine the sequence. As far as we know, this is the first study specifically focused on ensuring force safety in robotic manipulation. It significantly enhances the rationality of state planning, and the safe action trajectory is derived from inverse dynamics based on this refined planning. In practice, unlike previous approaches that concatenate visual and tactile data to generate future robot state sequences, our method employs tactile data as a calibration signal to adjust the robot’s state within the state space implicitly. Additionally, we’ve developed a large-scale simulation dataset called SafeDoorManip50k, offering extensive multimodal data to train and evaluate the proposed method. Extensive experiments show that our visual-tactile model substantially mitigates the risk of harmful forces in the door opening, across both simulated and real-world settings.
I Introduction
Vision-guided robotic manipulation [14, 7, 22, 13, 3, 2], which involves generating a feasible motion trajectory solely from visual information to carry out specific manipulation tasks, has become increasingly prevalent in the field of robotics and embodied AI. By leveraging visual inputs to perceive and interpret the environment, this approach allows robots to plan and execute intricate movements. However, current research predominantly emphasizes the success rate of these manipulation tasks, often neglecting the critical factor of how much force the robot should apply during the process. In practice, the inappropriate force exerted not only risks damaging the manipulated object but also puts unnecessary strain on the robot’s joints. Hence, addressing this issue is essential, and we refer to this consideration as ensuring force safety.
@@ I Introduction
Figure 1: The restoring force exerted by the robot’s end-effector can be decomposed into three components: $F_{x}$, $F_{y}$, and $F_{z}$. The component $F_{z}$ is tangent with the door’s opening trajectory and is termed the effective force. The forces lying in the xOy plane are orthogonal to the trajectory. These forces might cause damage to both the robot and the door and are therefore referred to as harmful forces.
@@
From the perspective of state planning, inappropriate force mainly arises when the generated state fails to meet the specific physical properties of the manipulated object. Taking the door-opening task shown in Fig. 1 as an example, the door can only move along the arc-shaped trajectory determined by its physical properties, such as size, opening angle, and its position relative to the robot. This indicates that all states of the robot’s end-effector, generated by the state planning model, must strictly adhere to this arc-shaped trajectory to ensure force safety. Otherwise, the robot controller might exert harmful forces in an attempt to reach states outside this trajectory, which could result in damage to both the door and the robot. To be brief, we define the state that lies on this arc-shaped trajectory as safe state, while those outside are deemed unsafe. In this paper, our primary focus is on planning safe states to ensure force safety throughout the robotic manipulation process.
@@
An intuitive solution for the above safe door-opening state planning can be found in bionics: when opening the door, humans estimate future door-opening states based on the door’s physical properties—such as size, opening angle, and so forth—using visual perception, and then modify them in real-time based on the forces sensed through tactile feedback during the actual door-opening process. Inspired by this, we aim to dynamically integrate real-time tactile feedback to refine the vision-guided generated states. However, this solution remains challenging due to the intricate, nonlinear dynamics between the current force feedback and the refinement of future states. These dynamics are influenced by factors like the robot’s manipulability, positions relative to the door, and other physical considerations in the real world.
To address such an issue, we develop a diffusion-based model named SafeDiff to plan safe states, leveraging the effectiveness of diffusion models in approximating complex distributions. In this work, we utilize offline demonstrations collected from the simulator to learn the aforementioned dynamics of the door opening and embed this knowledge into the state representation. This allows us to perform implicit calibration on vision-guided states online, utilizing real-time tactile feedback obtained during inference. Such a process enables the generated states to progressively satisfy the constraints imposed by the door’s properties, thereby ensuring force safety during the entire door-opening process.
Based on this, our SafeDiff demonstrates strong generalization capabilities and brings several benefits as well. (1) Robustness to External Disturbances. SafeDiff excels in handling environmental disturbances online. It can continuously correct the state changes induced by these disturbances during the manipulation. It ensures force safety, effectively adapting to the changing environmental conditions, while other methods often fail to open the door, let alone maintain force safety. This highlights SafeDiff’s ability to calibrate its planned state dynamically based on real-time tactile feedback. (2) Few-shot Sim-to-Real Transfer. SafeDiff demonstrates exceptional capability in transferring knowledge from simulation to the real world, even with limited real-world data. This significantly reduces the need for extensive real-world training, streamlining the transition from simulation to practical applications.
@@
The main contributions of this work are three-fold. 1) We introduce a novel benchmark centered on force safety during robotic manipulation. This benchmark includes three physically sound and computationally efficient metrics that can effectively assess the safety of state planning models. For the door opening, we create a large-scale safety-related simulation dataset, SafeDoorManip50k, to validate the effectiveness of the involved models. 2) We propose a diffusion-based model, SafeDiff, which dynamically incorporates real-time tactile feedback to calibrate vision-guided robot states implicitly. 3) Extensive experiments cross both simulated and real-world settings show that our method outperforms existing approaches in safe state planning, greatly reducing the risk of object damage during robotic manipulation.
II Related Works
II-A Vision-based Robotic Manipulation
@@
Numerous studies on vision-based robotic manipulation have addressed tasks such as object grasping [14, 7], articulated object manipulation [22, 13], and object reorientation [3]. These works emphasize improving the robot’s environmental perception through various visual modalities to enhance task success rates. For instance, [14, 7, 2] proposed using RGB-only images for robust robotic manipulation, while SAGCI [20], RLAfford [13], and Flowbot3D [9] rely solely on point clouds for observations. Additionally, [32, 29] integrated both RGB images and point clouds to promote the performance on specific manipulation tasks. However, the objects manipulated by robots are often fragile, especially articulated ones. In view of this, vision-based manipulation is challenging to apply in real-world applications because it cannot accurately reflect the force safety status of the manipulated objects. Therefore, it is of great significance for robots to incorporate tactile feedback such that it can dynamically adjust the planned states and handle objects in a safer and more controllable manner.
II-B Multimodal Tactile Feedback for Enhanced Manipulation
@@
Various learning-based approaches have employed tactile feedback to enhance robotic manipulation. For instance, [23] introduced a tactile perception-driven method that enables robots to learn how to grasp objects without relying on visual input. Numerous studies focus on grasp stability [8, 6, 4], as well as regrasping [5, 25]. A few methods [27, 15, 26, 28] combine reinforcement learning with tactile feedback to formulate manipulation strategies. And very few approaches leverage the combined benefits of both vision and touch. For example, [10] integrated prior knowledge with dynamic model adaptation to locally compensate for changing dynamics, while [16] developed a self-supervised learning framework that fuses visual and tactile inputs for peg insertion, improving learning efficiency. However, the majority of these works used tactile feedback to improve manipulation effectiveness rather than to guide safe planning.
Figure 2: Our framework takes a noise sequence as input, visual information, current robot state, and its corresponding force feedback as conditions and outputs the final safe states through $T$ denoising iterations. The architecture consists of an encoder and a decoder. The encoder is composed of a series of multi-scale Vision-Guided Mapping Modules (VMMs) that integrate visual data using FiLM [24] and generate state representations initially. The decoder comprises a stack of Tactile-Guided Calibration Modules (TCMs) which can refine the state representations based on tactile feedback.
II-C Datasets for Door Opening
@@
In recent years, a primary approach for door manipulation tasks has been to build simulation environments that emulate real-world conditions. Studies such as [9, 11, 12, 21, 30, 32, 33] have introduced a variety of simulated door-opening mechanisms, including pushing, pulling, and even those involving latching mechanisms. Moreover, datasets like PartNet-Mobility [31] and AKB-48 [19] offer diverse collections of articulated objects, including doors, but their focus on visual data collection overlooks crucial modalities such as tactile information, limiting their effectiveness for safe door-opening states planning. To address these shortcomings, we developed a comprehensive door manipulation environment with multi-modal inputs and provided a large-scale door-opening dataset to support safe manipulation planning.
III Methodology
III-A Preliminary
@@
We begin by briefly reviewing the diffusion models, a class of generative models that synthesize data by reversing a Markovian process where Gaussian noise is progressively added to data samples. These models consist of two primary phases: the forward process and the reverse process. In the forward process, the original data is systematically corrupted, transitioning from a structured state to pure Gaussian noise over a predefined number of steps, described by the equation $x_{t}=\sqrt{\alpha_{t}}x_{t-1}+\sqrt{1-\alpha_{t}}\epsilon$, where $\epsilon$ is Gaussian noise and $\alpha_{t}$ are variance-preserving coefficients. The reverse process entails learning to undo the noise addition to recover the original data from its noisy state. This involves training a neural network to estimate the reverse conditional distribution $p(x_{t-1}|x_{t})$, utilizing advanced deep learning techniques. A typical application of the diffusion model in robotic manipulation is Decision Diffuser [1], which makes decisions using a return-conditional diffusion model, allowing policies to generate behaviors satisfying individual constraints.
III-B The Overall Framework
@@
Motivated by the Decision Diffuser [1], the proposed SafeDiff aims to generate a consistent robot state sequence $\mathbf{S}=\{S_{k}\}_{k=1}^{L}$ that ensures force safety conditioned on visual-tactile information experienced during manipulation, thereby preventing any potential damage to the door. As shown in Fig. 2, we employ an encoder-decoder architecture for our diffusion model. Given the visual representation $\bf{O}$ of the current scene context, typically obtained from the image $\bf{I}$, and the current robot state $\bf{R}$, the initial input to the model is a set of Gaussian noise ${\bf{N}}=\{N_{k}\}_{k=1}^{L}$ with length $L$. After $T$ iterations, the model produces a sequence of $L$ consecutive robot states $\mathbf{S}$. Notably, we have opted to replace the action sequence generated in [1] with a sequence of robot states. This option stems from the fact that, while the door opening trajectory is predictable, conventional control actions do not inherently guarantee force safety. Instead, each robot state is closely correlated with the current state’s potential harmful force magnitude. Consequently, using robot states facilitates a more robust and efficient model training when integrating tactile feedback.
@@
To harness visual and tactile information effectively to generate safe and reasonable robot states, we first introduce the Vision-Guided Mapping Module (VMM) to construct the encoder for our state diffusion model. This module translates the robot’s current state, denoted as $\hat{\bf{S}}$, and the visual scene context $\bf{O}$, including the door size and relative position to the robot, into a comprehensive state space representation. Although the diffusion model can initially estimate robot state trajectories based on these visual cues, it falls short of guaranteeing force safety during the manipulation process. To tackle this, we further introduce a Tactile-Guided Calibration Module (TCM) to act as the decoder of our model. Drawing inspiration from human adaptability in responding to tactile feedback and adjusting actions accordingly, this module is designed to capture the intricate, nonlinear dynamics between the current force feedback, represented by $\bf{F}$, and the projected residuals of future states. For more details about the module design, please refer to Sec. III-C.
TABLE I: Quantitative evaluation of our method and existing models on the simulation scenarios from our SafeDoorManip50k, highlighting the effectiveness of our method in safe state planning. Ours (V) denotes our method utilizing only visual data as input, while Ours (V+T) incorporates both visual data and tactile calibration. The symbols ✓and ✗ indicate whether the door manipulated is seen or unseen.
Seen (?) SuR (%) $\uparrow$ AHF (N) $\downarrow$ Threshold - 5 N Threshold - 10 N Threshold - 15 N
SaR-95 (%) $\uparrow$ SaR-80 (%) $\uparrow$ SaR-95 (%) $\uparrow$ SaR-80 (%) $\uparrow$ SaR-95 (%) $\uparrow$ SaR-80 (%) $\uparrow$
Li et al. [17] 96.40 7.68 0.10 0.66 6.09 43.12 56.85 94.19
Ours (V) 99.98 6.27 0.83 7.87 22.41 57.65 66.86 88.26
Ours (V+T) 99.98 5.00 6.10 25.28 49.25 78.73 78.79 93.00
Li et al. [17] 95.27 7.51 0.00 0.66 8.82 47.55 58.15 92.76
Ours (V) 99.78 13.08 0.87 2.57 8.15 21.35 33.44 51.75
Ours (V+T) 99.33 5.08 5.13 24.90 55.54 79.33 79.87 91.54
III-C Network Architecture Design
@@
Visual-Guided Mapping Module (VMM) As shown in Fig. 2, we stack a series of VMMs with different temporal scales to construct the encoder for our state diffusion model. In this module, we initially generate the robot state representation by using visual information and the current robot state. Firstly, we use a Multi-Layer Perceptron (MLP), followed by a Resnet block (Res), to extract current scene context from the input image $\mathbf{I}$ and current state $\mathbf{\hat{S}}$. And following FiLM [24], we regard such extracted current scene context as affine coefficients and map Gaussian noise inputs $\mathbf{N}$ into the initial state representation. Then, a self-attention (Sttn) and a Resnet block (Res) are used to enhance the temporal coherence of these state representations:
$[\alpha,~{}\beta]$ $=\texttt{MLP}(\mathbf{I},~{}~{}\mathbf{\hat{S}})$ (1)
@@
IV Dataset
To overcome the gap in datasets, we establish the first dataset for ensuring force safety in door opening manipulation planning, named SafeDoorManip50k. Drawing on the open-source assets detailed in [18], we constructed a diverse collection of $57$ doors, each featuring unique structural designs and distinct color textures. Notably, due to functional limitations of the Isaac Gym, x-axis harmful forces are inaccurate with the original door handle. Consequently, we made modifications to the collision mesh of the door handle model, enabling accurate readings of the harmful forces in the x-axis. These doors were subsequently divided into a set of $45$ seen doors and a set of $12$ unseen doors.

### v2

@@ Abstract
In unstructured environments, robotic manipulation tasks involving objects with constrained motion trajectories—such as door opening—often experience discrepancies between the robot’s vision-guided end-effector trajectory and the object’s constrained motion path. Such discrepancies generate unintended harmful forces, which, if exacerbated, may lead to task failure and potential damage to the manipulated objects or the robot itself. To address this issue, this paper introduces a novel diffusion framework, termed SafeDiff. Unlike conventional methods that sequentially fuse visual and tactile data to predict future robot states, our approach generates a prospective state sequence based on the current robot state and visual context observations, using real-time force feedback as a calibration signal. This implicitly adjusts the robot’s state within the state space, enhancing operational success rates and significantly reducing harmful forces during manipulation. Additionally, we develop a large-scale simulation dataset named SafeDoorManip50k, offering extensive multimodal data to train and evaluate the proposed method. Extensive experiments show that our visual-tactile model substantially mitigates the risk of harmful forces in the door opening task, across both simulated and real-world settings. Project page is available at this URL.
11footnotetext: Equal contribution. Corresponding author.
Keywords: robotic manipulation, force safety, diffusion models, door opening
1 Introduction
@@
Figure 1: The restoring force exerted by the robot’s end-effector can be decomposed into three components: $F_{x}$, $F_{y}$, and $F_{z}$. The component $F_{z}$ is tangent with the door’s opening trajectory and is termed the effective force. The forces lying in the xOy plane are orthogonal to the trajectory. These forces might cause damage to both the robot and the door and are referred to as harmful forces.
In industrial and everyday settings, robotic manipulation tasks often involve objects whose motion trajectories are inherently constrained, such as opening doors, closing windows, pulling drawers, or assembling bolts and pins. Under visual guidance, the motion trajectory generated for the robot end-effector may deviate from the constrained trajectory of the manipulated object, leading to unintended additional forces at the end-effector, as illustrated in Fig. 1. As these mismatches increase, they can cause task failure, and the resulting harmful forces may even damage the manipulated object or the robot’s joint motors. Therefore, ensuring precise force regulation is critical for both safe and efficient manipulation. In this work, we define that maintaining the harmful forces during manipulation remain within a safe threshold as force safety.
Traditional approaches to force safety primarily rely on impedance control, which regulates the stiffness and damping characteristics of the robot’s end-effector to adapt to external forces and achieve compliant motion. However, these methods necessitate explicit modeling of interaction dynamics, making them well-suited for structured environments where system parameters can be accurately defined. As the demand for robotic manipulation in unstructured settings grows, end-to-end deep learning approaches [li2022see, brohan2023rt1roboticstransformerrealworld] have emerged as a promising alternative, offering enhanced adaptability and data-driven force regulation without the need for explicit system modeling. Despite notable advancements, existing research predominantly focuses on improving task success rates, often neglecting the critical aspect of precise force control.
@@
To address this gap, we propose a novel deep learning-based state planning approach to enhance force safety in robotic manipulation. Unlike traditional force control methods, our approach employs pure position control, significantly reducing hardware requirements while maintaining adaptability across diverse unstructured manipulation tasks. From the perspective of state planning, force safety issues mainly arise when the generated state fails to meet the specific physical properties of the structured environments. Taking the door-opening task shown in Fig. 1 as an example, the door can only move along the arc-shaped trajectory determined by its physical properties (e.g. the door’s size, opening angle, and position relative to the robot). This indicates that all states of the robot’s end-effector, generated by the state planning model, must strictly adhere to this arc-shaped trajectory to ensure force safety. Otherwise, the robot controller will attempt to reach states outside this trajectory, resulting in harmful forces. To be brief, we define the state that lies on this arc-shaped trajectory as safe state, while those outside are deemed unsafe. In this paper, our primary focus is on planning safe states to ensure force safety throughout the robotic manipulation process.
@@
An intuitive solution for the above safe door-opening state planning can be found in bionics: when opening the door, humans estimate future door-opening states based on the door’s physical properties—such as size, opening angle, and so forth—using visual perception, and then modify them in real-time based on the forces sensed through tactile feedback during the actual door-opening process. Inspired by this, we aim to dynamically integrate real-time tactile feedback to refine the vision-guided generated states. However, this solution remains challenging due to the intricate, nonlinear dynamics between the current force feedback and the refinement of future states. These dynamics are influenced by factors like the robot’s manipulability, positions relative to the door, and other physical considerations in the real world. To address such an issue, we develop a diffusion-based model named SafeDiff to plan safe states, leveraging the effectiveness of diffusion models in approximating complex distributions. In this work, we utilize offline demonstrations collected from the simulator to learn the aforementioned dynamics of the door opening and embed this knowledge into the state representation. This allows us to perform implicit calibration on vision-guided states online, utilizing real-time tactile feedback obtained during inference. Such a process enables the generated states to progressively satisfy the constraints imposed by the door’s properties, thereby ensuring force safety during the entire door-opening process.
@@
This work makes three key contributions: (1) We propose SafeDiff, a diffusion-based model that integrates real-time tactile feedback to implicitly calibrate vision-guided robot states, achieving robustness against external disturbances and maintaining force safety where prior methods often fail. (2) We demonstrate that SafeDiff achieves superior safe state planning across both simulation and real-world experiments, with strong few-shot sim-to-real transfer that greatly reduces real-world data requirements and minimizes object damage risk. (3) We introduce a novel benchmark for force safety in robotic manipulation, including three physically grounded, computationally efficient metrics and the large-scale simulation dataset SafeDoorManip50k for door-opening tasks.
2 Related Works
2.1 Vision-based Robotic Manipulation
@@
Numerous studies on vision-based robotic manipulation have addressed tasks such as object grasping [ichnowski2021dex, dai2023graspnerf], articulated object manipulation [mo2019partnet, geng2023rlafford], and object reorientation [andrychowicz2020learning]. These works emphasize improving the robot’s environmental perception through various visual modalities to enhance task success rates. For instance, [ichnowski2021dex, dai2023graspnerf, an2024rgbmanip] proposed using RGB-only images for robust robotic manipulation, while SAGCI [lv2022sagci], RLAfford [geng2023rlafford], and Flowbot3D [eisner2022flowbot3d] rely solely on point clouds for observations. Additionally, [xu2022universal, wu2020grasp] integrated both RGB images and point clouds to promote the performance on specific manipulation tasks. However, the objects manipulated by robots are often fragile, especially articulated ones. In view of this, vision-based manipulation is challenging to apply in real-world applications because it cannot accurately reflect the force safety status of the manipulated objects. Therefore, it is of great significance for robots to incorporate tactile feedback such that it can dynamically adjust the planned states and handle objects in a safer manner.
2.2 Multimodal Tactile Feedback for Enhanced Manipulation
@@
Various learning-based approaches have employed tactile feedback to enhance robotic manipulation. For instance, [murali2018learning] introduced a tactile perception-driven method that enables robots to learn how to grasp objects without relying on visual input. Numerous studies focus on grasp stability [dang2012learning, cui2020self, bekiroglu2013probabilistic], as well as regrasping [calandra2018more, su2015force]. A few methods [van2016stable, kalakrishnan2011learning, sung2017learning, van2015learning] combine reinforcement learning with tactile feedback to formulate manipulation strategies. And very few approaches leverage the combined benefits of both vision and touch. For example, [fu2016one] integrated prior knowledge with dynamic model adaptation to locally compensate for changing dynamics, while [lee2019making] developed a self-supervised learning framework that fuses visual and tactile inputs for peg insertion, improving learning efficiency. However, the majority of these works used tactile feedback to improve manipulation effectiveness rather than to guide safe planning.
2.3 Datasets for Door Opening
@@
In recent years, a primary approach for door manipulation tasks has been to build simulation environments that emulate real-world conditions. Studies such as [eisner2022flowbot3d, geng2023partmanip, geng2023gapartnet, mo2021where2act, wu2021vat, xu2022universal, zhu2020robosuite] have introduced a variety of simulated door-opening mechanisms, including pushing, pulling, and even those involving latching mechanisms. Moreover, datasets like PartNet-Mobility [xiang2020sapien] and AKB-48 [liu2022akb] offer diverse collections of articulated objects, including doors, but their focus on visual data collection overlooks crucial modalities such as tactile information, limiting their effectiveness for safe door-opening states planning. To address these shortcomings, we developed a comprehensive door manipulation environment with multi-modal inputs and provided a large-scale door-opening dataset to support safe manipulation planning.
3 Methodology
3.1 Preliminary
@@
We begin by briefly reviewing the diffusion models, a class of generative models that synthesize data by reversing a Markovian process where Gaussian noise is progressively added to data samples. These models consist of two primary phases: the forward process and the reverse process. In the forward process, the original data is systematically corrupted, transitioning from a structured state to pure Gaussian noise over a predefined number of steps, described by the equation $x_{t}=\sqrt{\alpha_{t}}x_{t-1}+\sqrt{1-\alpha_{t}}\epsilon$, where $\epsilon$ is Gaussian noise and $\alpha_{t}$ are variance-preserving coefficients. The reverse process entails learning to undo the noise addition to recover the original data from its noisy state. This involves training a neural network to estimate the reverse conditional distribution $p(x_{t-1}|x_{t})$, utilizing advanced deep learning techniques.
Figure 2: Our framework takes a noise sequence as input, visual information, current robot state, and its corresponding force feedback as conditions and outputs the final safe states through $T$ denoising iterations. The architecture consists of an encoder and a decoder. The encoder is composed of a series of multi-scale Vision-Guided Mapping Modules (VMMs) that integrate visual data using FiLM [perez2018film] and generate state representations initially. The decoder comprises a stack of Tactile-Guided Calibration Modules (TCMs) which can refine the state representations based on tactile feedback.
A typical application of the diffusion model in robotic manipulation is Decision Diffuser [ajay2022conditional], which makes decisions using a return-conditional diffusion model, allowing policies to generate behaviors satisfying constraints.
3.2 The Overall Framework
@@
Motivated by the Decision Diffuser [ajay2022conditional], the proposed SafeDiff aims to generate a consistent robot state sequence $\mathbf{S}=\{S_{k}\}_{k=1}^{L}$ that ensures force safety conditioned on visual-tactile information experienced during manipulation, thereby preventing any potential damage to the door. As shown in Fig. 2, we employ an encoder-decoder architecture for our diffusion model. Given the visual representation $\bf{O}$ of the current scene context, typically obtained from the image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$, and the current robot state $\mathbf{R}\in\mathbb{R}^{7+6+7+7}$ (representing end effector pose, velocity, and joint position, velocity respectively), the initial input to the model is a set of Gaussian noise $\mathbf{N}\in\mathbb{R}^{L\times 7}$} with length $L$. After $T$ iterations, the model produces a sequence of $L$ consecutive robot states $\mathbf{S}\in\mathbb{R}^{L\times 7}$. Notably, we have opted to replace the action sequence generated in [ajay2022conditional] with a sequence of robot states. This option stems from the fact that, while the door opening trajectory is predictable, conventional control actions do not inherently guarantee force safety. Instead, each robot state is closely correlated with the current state’s potential harmful force magnitude. Consequently, using robot states facilitates a more robust and efficient model training when integrating tactile feedback.
@@
To harness visual and tactile information effectively to generate safe and reasonable robot states, we first introduce the Vision-Guided Mapping Module (VMM) to construct the encoder for our state diffusion model. This module translates the robot’s current state, denoted as $\hat{\bf{S}}$, and the visual scene context $\bf{O}$, including the door size and relative position to the robot, into a comprehensive state space representation. Although the diffusion model can initially estimate robot state trajectories based on these visual cues, it falls short of guaranteeing force safety during the manipulation process. To tackle this, we further introduce a Tactile-Guided Calibration Module (TCM) to act as the decoder of our model. Drawing inspiration from human adaptability in responding to tactile feedback and adjusting actions accordingly, this module is designed to capture the intricate, nonlinear dynamics between the current force feedback, represented by $\bf{F}$, and the projected residuals of future states. For more details about the module design, please refer to Sec. 3.3.
3.3 Network Architecture Design
@@
Visual-Guided Mapping Module (VMM) As shown in Fig. 2, we stack a series of VMMs with different temporal scales to construct the encoder for our state diffusion model. In this module, we initially generate the robot state representation by using visual information and the current robot state. Firstly, we use a Multi-Layer Perceptron (MLP), followed by a Resnet block (Res), to extract current scene context from the input image $\mathbf{I}$ and current state $\mathbf{\hat{S}}$. And following FiLM [perez2018film], we regard such extracted current scene context as affine coefficients and map Gaussian noise inputs $\mathbf{N}$ into the initial state representation. Then, a self-attention (Sttn) and a Resnet block (Res) are used to enhance the temporal coherence of these state representations:
[…]
@@
Implementation The proposed network follows a multi-scale architecture inspired by U-Net, enabling hierarchical feature extraction and reconstruction. During the encoding phase, VMM is employed for progressive downsampling, successively reducing the sequence length from L to L/2, L/4, L/8, and L/16, while correspondingly increasing the feature dimensions from 3 to 32, 64, 128, and 256. By extracting hierarchical visual features, the encoder captures multi-scale environmental information essential for accurate trajectory generation. In the decoding phase, TCM is utilized for progressive upsampling, integrating tactile features to refine the generated trajectory and enhance adaptability to environmental constraints. To facilitate effective gradient propagation across different scales, shortcut connections are incorporated between all corresponding encoder and decoder layers. This design enhances optimization stability and preserves fine-grained details.
4 Dataset
To overcome the gap in datasets, we establish the first dataset for ensuring force safety in door opening manipulation planning, named SafeDoorManip50k. Drawing on the open-source assets detailed in [li2024unidoormanip], we constructed a diverse collection of $57$ doors, each featuring unique structural designs and distinct color textures. Notably, due to functional limitations of the Isaac Gym, x-axis harmful forces are inaccurate with the original door handle. Consequently, we made modifications to the collision mesh of the door handle model, enabling accurate readings of the harmful forces in the x-axis. These doors were then divided into a set of $45$ seen doors and a set of $12$ unseen doors.